# Is ‘Right’ Right? Enhancing Object Orientation Understanding in Multimodal Language Models through Egocentric Instruction Tuning

### v1

@@ 2.2 Object Orientation Understanding
More recently, MMVP [43] has underscored the need for object orientation understanding within MLLMs. To evaluate visual encoder representations, MMVP introduces a simple evaluation method and dataset that measure MLLMs’ ability to interpret object orientation. However, this approach is limited by a small sample size of real-image data, which restricts comprehensive assessment across diverse image domains and various orientation complexities. To address this limitation, our study proposes a large-scale benchmark specifically designed to evaluate MLLMs’ object orientation understanding across a broad range of domains and conditions, facilitating a more robust analysis of MLLM capabilities.
@@ 4 EgoOrientBench: Egocentric Orientation Understanding Benchmark
MLLMs, known for their versatile capabilities, can be applied across various domains. They are valuable not only in realistic image contexts [32, 41] but also in applications like AI docents for artwork [25] and object recognition in augmented reality environments [38]. To evaluate MLLMs’ ability to recognize object orientation across diverse settings, we construct a large-scale benchmark dataset, EgoOrientBench, for orientation understanding from a user-centered, egocentric perspective, using data collected from five distinct image datasets:
@@ 5.1 Experimental Setup
Benchmarks. To verify the effectiveness of egocentric instruction tuning, we conduct experiments in EgoOrientBench to assess MLLMs’ understanding of object orientation. Additionally, to determine whether egocentric instruction tuning enhances object orientation understanding without compromising general response generation, we evaluate performance on the MME benchmark [15], which measures MLLMs’ overall response generation capabilities.
@@ 5.2 Experimental Result
Method Category Total
Backbone Method Perception Cognition
LLaVA 1.5 Zero-shot 1476.2 316.6 1792.8
Ours 1453.8 299.0 1752.8
mPLUG-Owl2 Zero-shot 1384.3 322.1 1706.3
Ours 1401.2 326.1 1727.3
InternVL2-4B Zero-shot 1547.6 541.1 2088.7
Ours 1514.7 531.2 2045.9
Table 2: Evaluation results on the MME benchmark. While understanding of object orientation improves through egocentric instruction tuning, the MME evaluation results show only slight degradation or, in the case of mPLUG-Owl2, even improvement.
Results on MME. Table 2 presents the experimental results on the MME benchmark. The results indicate that performance remains stable after applying our egocentric instruction tuning, with no significant degradation compared to pre-tuning levels. Notably, mPLUG-Owl2 shows improved performance in the Perception and Cognition categories. These findings demonstrate that our instruction tuning effectively aligns MLLM’s object orientation understanding with the user’s perspective while maintaining the model’s overall response generation capabilities.
@@ 6.1 Applicability in Real-world Scenarios
As shown in the Table 4, the experimental results demonstrate that our egocentric instruction tuning effectively improves spatial reasoning abilities for understanding preposition relationships between objects. Specifically, accuracy improved by 15.0%p, 14.3%p, and 3.4%p in LLaVA-1.5, mPLUG-Owl2, and InternVL2, respectively. This indicates that our approach can enhance MLLM performance in user-centered applications that require recognizing object positions, understanding orientations, and interpreting inter-object relationships.
@@ 6.2 Limitations
Figure 7 shows examples of error cases for LLaVA-1.5 after applying our egocentric instruction tuning. In the image, the zebra and giraffe have different orientations for their bodies and heads. Unlike inanimate objects like cars or tractors, living objects such as zebras and humans can exhibit these complex, independent orientations. These complexities challenge our Response Type 2 data, which connects object detail recognition with orientation understanding—for instance, when the head and tail imply different orientations. They also complicate Response Type 3 data, which relies on interpreting orientation relationships for alignment tasks. In future work, we plan to address these challenges by moving beyond the current discrete orientation categories.
@@ 7 Conclusion
In this study, we identify inconsistent object orientation annotations as a primary factor hindering MLLM’s orientation understanding and propose an Egocentric Instruction Tuning method to address this by aligning MLLM’s object orientation comprehension with the user’s egocentric perspective. Additionally, we introduce EgoOrientBench, a benchmark designed to comprehensively evaluate MLLM’s orientation understanding across three tasks and five image datasets. Experimental results on this benchmark, as well as on the MME benchmark, demonstrate that our egocentric instruction tuning effectively enhances object orientation understanding without compromising general response generation capabilities. Our analysis also reveals that current MLLMs exhibit a strong bias toward specific orientations, which our instruction tuning mitigates. Finally, we apply the enhanced MLLM, trained with egocentric instruction tuning, to pedestrian walking direction prediction and spatial reasoning tasks involving spatial relationships between objects. This demonstrates the practical benefits of enhanced orientation understanding for real-world applications. We believe our instruction tuning method and benchmark will significantly contribute to future research on MLLM’s object orientation understanding.

### v2

# Is ‘Right’ Right? Enhancing Object Orientation Understanding in Multimodal Large Language Models through Egocentric Instruction Tuning

@@ 2.2 Object Orientation Understanding
More recently, MMVP [42] has underscored the need for object orientation understanding within MLLMs. To evaluate encoder representations, MMVP introduces a simple evaluation method and dataset that measure MLLMs’ ability to interpret object orientation. However, this approach is limited by a small sample size of real-image data, which restricts comprehensive assessment across diverse image domains and various orientation complexities. To address this limitation, we propose a large-scale benchmark specifically designed to evaluate MLLMs’ object orientation understanding across a broad range of domains and conditions, facilitating a more robust analysis of MLLM capabilities.
@@ 4 EgoOrientBench: Egocentric Orientation Understanding Benchmark
MLLMs, known for their versatile capabilities, can be applied across various domains. To assess MLLMs’ ability to recognize object orientation in diverse domains, we introduce EgoOrientBench, a large-scale benchmark dataset designed for orientation understanding from a user-centered, egocentric perspective. The dataset is constructed from five distinct image datasets, spanning various domains such as drawings, paintings (from PACS and DomainNet), and 3D-rendered images (from OmniObject3D).
@@ 5.1 Experimental Setup
Benchmarks. To verify the effectiveness of egocentric instruction tuning, we conduct experiments in EgoOrientBench to assess MLLMs’ understanding of object orientation. Additionally, to determine whether egocentric instruction tuning enhances object orientation understanding without compromising general response generation, we evaluate the models’ generalization capabilities using four benchmarks: MME [16], MMStar [6], MMMU [50], and POPE [28].
@@ 5.2 Experimental Result
Method MME MMStar MMMU POPE
Backbone Method
LLaVA 1.5 Zero-shot 1792.8 34.67 35.11 82.03
Ours 1752.8 35.87 34.44 88.36
mPLUG-Owl2 Zero-shot 1706.3 34.33 37.55 86.16
Ours 1727.3 35.27 38.55 85.60
InternVL2-4B Zero-shot 2088.7 54.26 47.22 85.91
Ours 2045.9 53.13 48.00 85.56
Table 2: Experimental results evaluating the general performance of MLLMs on the MME [16], MMStar [6], MMMU [50], and POPE [28] before and after egocentric instruction tuning.
Results on general benchmarks. Table 2 presents the experimental results on the general performance benchmarks. The results indicate that performance remains stable after applying our egocentric instruction tuning, with no significant degradation compared to pre-tuning levels. Overall, the three models exhibit largely comparable performance across the four benchmarks. In particular, LLaVA 1.5 shows a significant improvement in POPE, while mPLUG-Owl2 demonstrates considerable gains in MME. Similarly, InternVL2 maintains stable or slightly improved performance without noticeable degradation across most benchmarks.
These findings demonstrate that our instruction tuning effectively aligns MLLM’s object orientation understanding with the user’s perspective while maintaining the model’s overall response generation capabilities.
@@ 6.1 Applicability in Real-world Scenarios
As shown in the Table 4, our egocentric instruction tuning effectively improves spatial reasoning abilities for understanding preposition relationships between objects. Specifically, accuracy improved by 15.0%p, 14.3%p, and 3.4%p in LLaVA-1.5, mPLUG-Owl2, and InternVL2, respectively. This indicates that our approach can enhance MLLM performance in user-centered applications that require recognizing object positions, understanding orientations, and interpreting their spatial relationships.
@@ 6.2 Limitations
Figure 7 presents examples of error cases in LLaVA-1.5 after applying our egocentric instruction tuning. In these examples, the zebra and giraffe exhibit different orientations for their bodies and heads. Unlike inanimate objects such as cars or tractors, living beings like zebras and humans can adopt complex, independent orientations. These complexities pose challenges for our Response Type 2 data, which links object detail recognition with orientation understanding—for instance, when the head and tail suggest different orientations. Similarly, they complicate Response Type 3 data, which depends on interpreting orientation relationships for alignment tasks. To address these challenges, future work will explore approaches beyond the current discrete orientation categories.
@@ 7 Conclusion
In this study, we identify inconsistent object orientation annotations as a primary factor hindering MLLM’s orientation understanding and propose an Egocentric Instruction Tuning method to address this by aligning MLLM’s object orientation comprehension with the user’s egocentric perspective. Additionally, we introduce EgoOrientBench, a benchmark designed to comprehensively evaluate MLLM’s orientation understanding across three tasks and five image datasets. Experimental results on this benchmark, as well as on the general performance benchmarks, demonstrate that our egocentric instruction tuning effectively enhances object orientation understanding without compromising general response generation capabilities. Our analysis also reveals that current MLLMs exhibit a strong bias toward specific orientations, which our instruction tuning mitigates. Finally, we apply the enhanced MLLM, trained with egocentric instruction tuning, to pedestrian walking direction prediction and spatial reasoning tasks involving spatial relationships between objects. This demonstrates the practical benefits of enhanced orientation understanding for real-world applications. We believe our instruction tuning method and benchmark will significantly contribute to future research on MLLM’s object orientation understanding.
Acknowledgment
This work was partly supported by the Institute of Information & Communications Technology Planning & Evaluation(IITP)-ICT Creative Consilience Program grant funded by the Korea government(MSIT)(IITP-2025-RS-2020-II201819) and the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (RS-2023-00212828, RS-2024-00350430).